OpenStudy (anonymous):

@iGreen i will give a meadel and fan please help jerom has 1/2 of the groups video games at his house. mario has 2/5 of the groups game at his house. what fraction of the groups video games is their at jeroms haouse or mario house. Give an example and explain.

OpenStudy (igreen):

Jerom's house OR Mario's house. or is it Jerom's house AND Mario's house.

OpenStudy (anonymous):

jerom and mori house @iGreen

OpenStudy (anonymous):

You have to make the two fractions the same. The easiest way to do this is by multiplying 2 * 5. now you have 5/10 at Jerom's house, and 4/10 at Mario's house.

OpenStudy (igreen):

Okay, just add the fractions. \(\dfrac{1}{2} + \dfrac{2}{5}\) Find the common denominator, in this case it's 10. So let's change the fractions to fractions with denominators of 10. \(\dfrac{1 \times 5}{2 \times 5} + \dfrac{2 \times 2}{5 \times 2}\) \(\dfrac{5}{10} + \dfrac{4}{10}\) Now just add the two fractions. \(\dfrac{5 + 4}{10}\) Can you add 5 + 4?

OpenStudy (anonymous):

9

OpenStudy (anonymous):

9

OpenStudy (igreen):

Yep, so your answer is \(\dfrac{9}{10}\).

OpenStudy (anonymous):

thanks i green

OpenStudy (igreen):

No problem.